# LittleBit: Ultra Low-Bit Quantization via Latent Factorization

### v3

@@ A.2.2 SmoothSign versus Straight-Through Estimator
For backpropagating through the non-differentiable $\mathrm{sign}(x)$ function, we compared the Straight-Through Estimator (STE) [ste2013] with our proposed SmoothSign technique. SmoothSign employs the $\mathrm{sign}(x)$ function for the forward pass. For the backward pass, it utilizes a smooth proxy gradient, specifically the derivative of $\tanh(kx)$ where $k=100$ (illustrated in Fig.˜8).
@@ A.2.3 Grouped Query Attention
Modern LLMs often use Grouped Query Attention (GQA) [gqa2023] or Multi-Query Attention (MQA) [fast2019], where key (K) and value (V) projection layers have smaller output dimensions than query (Q) projections. Applying LittleBit at ultra-low bits (e.g., 0.1 BPW) to these GQA/MQA models can result in an extremely small latent dimension $r$ for K/V layers (e.g., $r\approx 20$ for Llama3-8B K/V layers at 0.1 BPW overall), potentially creating an information bottleneck and hindering QAT or performance.
@@ A.3 Perplexity on C4 and PTB
Table 8: Perplexity (PPL) comparison on C4 and PTB validation sets for Llama2 models. Lower PPL indicates better performance. STBLLM [stbllm2024] utilizes N:M sparsity (ratio in parentheses).
@@ A.5 Pruning versus SVD
To select a base compression method to achieve sub-0.5 BPW, we compared pruning (Wanda [wanda2023]) with SVD-based compression (SVD-LLMv2 [svdllmv2_2025]) on Llama-7B ( Table˜9). SVD showed superior robustness at extreme compression ratios (e.g., 25% parameter retention).
@@
Remaining Parameter Ratio Pruning [wanda2023] SVD [svdllmv2_2025]

### v4

@@ A.2.2 SmoothSign versus Straight-Through Estimator
For backpropagating through the non-differentiable $\mathrm{sign}(x)$ function, we compared the Straight-Through Estimator (STE) [3] with our proposed SmoothSign technique. SmoothSign employs the $\mathrm{sign}(x)$ function for the forward pass. For the backward pass, it utilizes a smooth proxy gradient, specifically the derivative of $\tanh(kx)$ where $k=100$ (illustrated in Fig.˜8).
@@ A.2.3 Grouped Query Attention
Modern LLMs often use Grouped Query Attention (GQA) [2] or Multi-Query Attention (MQA) [47], where key (K) and value (V) projection layers have smaller output dimensions than query (Q) projections. Applying LittleBit at ultra-low bits (e.g., 0.1 BPW) to these GQA/MQA models can result in an extremely small latent dimension $r$ for K/V layers (e.g., $r\approx 20$ for Llama3-8B K/V layers at 0.1 BPW overall), potentially creating an information bottleneck and hindering QAT or performance.
@@ A.3 Perplexity on C4 and PTB
Table 8: Perplexity (PPL) comparison on C4 and PTB validation sets for Llama2 models. Lower PPL indicates better performance. STBLLM [10] utilizes N:M sparsity (ratio in parentheses).
@@ A.5 Pruning versus SVD
To select a base compression method to achieve sub-0.5 BPW, we compared pruning (Wanda [48]) with SVD-based compression (SVD-LLMv2 [55]) on Llama-7B ( Table˜9). SVD showed superior robustness at extreme compression ratios (e.g., 25% parameter retention).
@@
Remaining Parameter Ratio Pruning [48] SVD [55]